### v1

# CONLINE: Complex Code Generation and Refinement with Online Searching and Correctness Testing

Abstract
Large Language Models (LLMs) have revolutionized code generation ability by converting natural language descriptions into executable code. However, generating complex code within real-world scenarios remains challenging due to intricate structures, subtle bugs, understanding of advanced data types, and lack of supplementary contents. To address these challenges, we introduce the CONLINE framework, which enhances code generation by incorporating planned online searches for information retrieval and automated correctness testing for iterative refinement. CONLINE also serializes the complex inputs and outputs to improve comprehension and generate test case to ensure the framework’s adaptability for real-world applications. CONLINE is validated through rigorous experiments on the DS-1000 and ClassEval datasets. It shows that CONLINE substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality and reliability of LLMs in generating intricate code.
CONLINE: Complex Code Generation and Refinement
with Online Searching and Correctness Testing
Xinyi He1thanks: The contributions by Xinyi He, Jiaru Zou and Yun Lin have been conducted and completed during their internships at Microsoft.   Jiaru Zou2*   Yun Lin3*   Mengyu Zhou4thanks: Corresponding author.   Shi Han 4   Zejian Yuan1   Dongmei Zhang4 1 Xi’an Jiaotong University 2 University of Illinois at Urbana-Champaign 3 Peking University 4 Microsoft Research hxyhxy@stu.xjtu.edu.cn, linyun@stu.pku.edu.cn, jiaruz2@illinois.edu, yuan.ze.jian@xjtu.edu.cn, {mezho, shihan, dongmeiz}@microsoft.com
1 Introduction
Figure 1: An Example of the Human Programmer Code-writing Process Imitated by the CONLINE. After the problem is received, an online search is performed to simulate search results and create an initial version of code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
Automated code generation from natural language descriptions has became more realistic, as large language models (LLMs) show their potential to generate complex and accurate code Manna and Waldinger (1971); Li et al. (2023); Luo et al. (2023); Rozière et al. (2024). There are a variety of methods that have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al. (2023); Jiang et al. (2023) and debugging generated code Zhang et al. (2023); Chen et al. (2023).
In real-world scenarios, complex code generation can involve intricate problem description, sophisticated code logic, and advanced data types Lai et al. (2022); Du et al. (2023); He et al. (2023). For example (see Figure 1), solving the required code generation involves multiple transformations of various matrices, as well as the invocation of functions that demands numerous and stringent parameters. Existing methods were not enough to address the following new challenges that emerge during complex code generation:
  1. 1.
    Complex code generation is hindered by intricate structures and logic. E.g., the complexity of the example in Figure 1 is as previously described.
  2. 2.
    Obscure bugs in complex code necessitate critical detection and refinement. Existing methods often refine code based on execution errors testing Zhang et al. (2023); Jiang et al. (2023), but for complex codes, sometimes the executable code has non-obvious bugs,e.g., initial code in Figure 1.
  3. 3.
    The complexity of inputs and outputs hampers LLM’s comprehension ability during code refinement. In complex code, inputs and outputs like Pandas DataFrames, PyTorch tensors, and matplotlib PNG images are intricate. Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities and so on.
  4. 4.
    Real-world scenarios frequently lack supplemental content, such as test cases and documentation pools. Existing works predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023) or off-line document pools Zhou et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios.
To address the above challenges, we propose a new code generation framework CONLINE that enhances the code generation and refinement quality of LLMs with planned online searching and automated correctness testing steps in this paper. The intuition of CONLINE is straightforward: During the process of coding, most human programmers are not bothered by the above challenges as shown in Figure 1. They could easily bypass them by searching on-line for answers, experiences, or guidelines. Also, they are able to design test cases and run code against them to verify the correctness of the code logic.
Correspondingly, CONLINE uses online search for code generation to partially address Challenges 1 and 4. For Challenge 1, it offers following advantages: (1) Retrieving information from library documents allows the retrieval of standardized function syntax, preventing errors from model hallucinations e.g., generating non-existing parameters. (2) Retrieving information from the latest blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. For Challenge 4, online search encompasses the content of offline document pools and extends beyond with its superior capability to adapt to the diverse and evolving needs of real-world scenarios. Directly searching for solutions to such problems is difficult, because the problems are intricate and may involve multiple challenges. Therefore, we propose online search with generation query through planning.
To address the Challenge 2, this work proposes correctness testing in refinement. LLMs may generate complex code that contains subtle errors, resulting in code that is executable but does not produce the desired output. Therefore, during the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors. We propose incorporating both the execution output results and the errors as integral parts of the refinement to enhance the correctness of the code.
In refinement with correctness testing, we propose generation of test cases to address the Challenge 3. Test cases are crucial, as they serve as indispensable inputs for the code execution. This work utilizes LLMs to automatically generate test cases, making the framework applicable even in real-world scenarios where no known test cases exist. To address the Challenge 4, we propose serialization of input and output. It renders the input and output more intuitive and understandable for the model, particularly for those that are excessively long or non-textual modalities.
We test the effectiveness of CONLINE on two complex code generation datasets using three different LLMs. For the complex code generation datasets, we selected DS-1000, which focuses on complex data science code generation, and the Class Eval dataset, which is oriented towards class-level code generation. By comparing with the existing state-of-the-art (SOTA) baseline, we achieved a 10.9% improvement on DS-1000 and average 9.47% on ClassEval, confirming the effectiveness of CONLINE. Through comprehensive ablation studies, we verified the effectiveness of each component of CONLINE. Moreover, given the varying performance of CONLINE on different base models, we discovered that CONLINE requires models to have capabilities such as Planning, which vary according to the complexity of the problem, in order to further enhance the model’s code generation abilities.
2 Related Work
The realm of automated code generation has been propelled by benchmark datasets such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021), which assess the proficiency of language models in generating executable code from descriptions. These datasets encompass a variety of programming problems, yet recent studies have sought to escalate the complexity of code generation tasks. Works like DS-1000 Lai et al. (2022), ClassEval Du et al. (2023) and Text2Analysis He et al. (2023) have introduced datasets targeting specialized domains, including data science, object-oriented class generation, and data analysis. These endeavors reflect an emerging trend towards enhancing models’ abilities to produce sophisticated and domain-specific code structures.
Recent research has concentrated on improving Large Language Models (LLMs) for code generation via retrieve, refine and so on strategies. Retrieve enhances model performance by pulling in external information, with techniques like DocPrompt Zhou et al. (2023) accessing fixed document repositories and SELVEVOLVE Jiang et al. (2023) using models as knowledge bases. However, these are limited to document-sourced content and provide limited benefits.
Refine iteratively enhances generated code for greater precision. Projects such as Self-Debug Chen et al. (2023) and SELFEVOLVE Jiang et al. (2023) focus on error refinement, while Self-EditZhang et al. (2023) uses code execution results as feedback. Despite advancements, these methods often miss subtle bugs and are less applicable to complex, real-world code generation without readily available test cases or ground truths.
Figure 2: The Pipline of CONLINE. Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps. These queries enable the retrieval of diverse information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases for testing the initial code. The test results serve as crucial inputs for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
3 Methodology
The code generation task involves predicting a solution code $W$ given a problem description $D$. When given an input $i$, the execution of code $W$ produces an output result $o$ and a potential error $e$, where both $o$ and $e$ can be empty $\varnothing$. The generated codes are evaluated against a set of test cases and ground truth $\{(t_{j},g_{j})\}_{j=1}^{J}$. The correctness of the code $W$ is determined by verifying $o_{j}=g_{j}\land e_{j}=\varnothing$ when all $i_{j}=t_{j},j\in\{1,\ldots,J\}$.
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieve, where relevant information is obtained through online saerch and utilized by LLMs to generate initial code. The second step is refine, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
3.1 Retrieve
The difficulty in achieving effective online retrieval lies in formulating optimal search queries. On the one hand, for complex code generation, the problems are intricate and may involve multiple challenges. Directly searching for solutions to such problems is inaccurate and difficult. On the other hand, it is challenging that matching queries directly through methods for offline document like similarity calculations, due to the nature of online libraries. Therefore, we propose using planning to solve the challenge.
The retrieval process is divided into three steps: 1. Search queries $Q=\{q_{1},\ldots,q_{N}\}$ are generated through planning. 2. Conducting online searches using these queries to obtain relevant background information $INFO=\{info_{1},\ldots,info_{M}\}$. 3. The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $I$:
$\widehat{W}_{0}\sim p_{\theta}(.|D,I)$ (1)
3.1.1 Generation Query through Planning
In order to generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem. The planning phase involves outlining the natural language steps $P=\{plan_{1},\ldots,plan_{N}\}$ required to address the problem. Later, the assessment involves utilizing LLMs to determine whether each planning step requires an online search. Subsequently, the planning steps identified as necessitating online search are translated into queries $Q=\{q_{1},\ldots,q_{N}\}$ for use in the subsequent search process.
$\widehat{P},\widehat{Q}\ \sim p_{\theta}(.|D)$ (2)
3.1.2 Online Search
For the above generated queries, we conduct an online search. We use the online search API [^1] https://github.com/Nv7-GitHub/googlesearch for the search process.
$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$
where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ and $N_{u}=1$.
Through the analysis of the website distribution Table 4, we observed that more than 90% of the URLs are concentrated on a total of 8 websites. Specific extraction rules are established for prominent websites such as StackOverflow to extract key information, facilitating a more comprehensive understanding of the website’s content by subsequent models. Generic extraction rules are employed for extracting key information from other websites.
$info_{j,k}=extract(url_{k}),k\in\{1,\ldots,N_{u}\}$
The information $INFO$ is composed of details from each query $q_{j}$, each url $url_{k}$, with each piece of information $info_{j,k}$ extracted.
3.2 Refine
Existing work Chen et al. (2023); Jiang et al. (2023) typically emphasizes correction of errors identified during the refinement process. However, we observe that refining code that produces error-free outputs is equally crucial during the refinement process. Therefore, we introduce correction testing in §3.2.1. Additionally, we propose methods for the generation of test cases and serialization of inputs and outputs during the refinement process.
3.2.1 Correctness Testing
Correctness testing refers to the refinement of generated code based on correctness, determined by analyzing errors and output results obtained during code execution. In the context of complex code generation, the intricate logic of the code makes it challenging for the LLMs to consider every detail during code generation, and precisely ascertain the results obtained at each step of the execution process. Consequently, some code may be executed without errors, producing output results that do not align with what is expected. Incorporating both the error and the output result into the refinement process allows the model to take advantage of self-correction mechanisms.
$\left\{\begin{array}[]{ccl}e_{j,k},o_{j,k}&=&execute(W_{j},i_{k}),j\in\{1,\ldots,N_{f}\}\\ INFO_{e_{j,k}}&=&\{e_{j,k},extract(search(e_{j,k}))\}\\ \widehat{W}_{j+1}&\sim&p_{\theta}(.|D,W_{j},\{S_{i},S_{o_{j}},INFO_{e_{j}}\}_{k}),\\ &&\hfil k\in\{1,\ldots,N_{i}\}\end{array}\right.$
where, $N_{f}$ is the total number of refinement steps, $N_{i}$ is the number of inputs. $i_{K}$ is the k-th input for the problem from Equation (3), $S_{i}$ and $S_{o_{j}}$ is the serialization of the input and output from Equation (4).
3.2.2 Generation of Test Case
Test cases is crucial, as they serve as indispensable inputs for the code execution in refinement. While, existing works in refining code predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios. Moreover, some existing work Chen et al. (2023) even uses the ground truth output of the test case to refine the code, which is even more challenging to obtain for complex code problems in real-world scenarios. Because their problems involve various logical operations, and even with input test cases, deriving answers directly without code-based computations is demanding.
In this study, we propose generation of test case with LLMs to adapt to real-world scenarios.
$\left\{\begin{array}[]{ccl}\widehat{I}&\sim&p_{\theta}(.|D)\\ I&=&\{i_{1},\ldots,i_{N_{i}}\}\end{array}\right.$ (3)
3.2.3 Serialization of Input and Output
Serialization of input and output makes them more intuitive and understandable for the model. For complex code, some inputs and outputs are intricate, such as Pandas DataFrames, PyTorch tensors, and matplotlib PNG images. Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities, and so on.
In this study, we serialize common data structures in Python as follows:
  1. 1.
    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes a truncated string of data, data type, data shape, and statistical information.
  2. 2.
    For image structures (such as PNG images generated by the Matplotlib library), we serialize them into SVG (Scalable Vector Graphics) format for LLMs to comprehend.
$S_{n}=serialize(n),n\in\{i_{k},o_{j,k}\}$ (4)
Table 1: Table of Main Results and Ablation Study for DS-1000. The base model for CONLINE is GPT-4. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
Method Perturbation Total/Avg.
Origin Surface Semantic Diff-Rewrite
DocPrompting 53.95 50.00 38.39 21.05 43.30
Self-Debugging 63.38 59.21 45.65 28.40 53.00
SELFEVOLVE 66.23 67.11 48.70 33.95 57.10
CONLINE 71.71 74.34 66.96 53.09 68.00
   w/o refinement of output 68.42 69.74 62.61 48.77 64.10
   w/o refinement of error 68.20 73.03 62.61 49.38 64.60
   w/o serialization 70.18 75.00 65.22 51.23 66.70
   w/o generation of test case 66.23 71.05 59.57 45.68 62.10
   w/o online retrieval 68.64 70.39 60.00 51.23 64.10
   GPT-4 64.47 69.74 56.96 43.83 60.20
4 Experiment
4.1 Experiment Setup
4.1.1 Datasets
We conduct experience on two complex code generation datasets:
1. DS-1000 Lai et al. (2022): DS-1000 is a code generation benchmark with a thousand data science questions spanning seven Python libraries. The complexity of this dataset is manifested in two aspects. First, complexity arises from intricate logical reasoning required during code generation due to the complex nature of the problems. For example, on the DS-1000 dataset, the average length of problem descriptions is 140 words, whereas other commonly used code generation datasets such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) have lengths of 23 and 15.7 words, respectively. Secondly, the input-output involves various complex data structures related to data science, making the code logic intricate during transformations of the data.
The dataset provides both Insertion and Completion style prompts, where the data is the same, differing only in prompt format, thus yielding similar results. In this paper, experiments are conducted with the Completion style prompt.
2. ClassEval Du et al. (2023): ClassEval is the first class-level Python code generation benchmark designed to evaluate code generation models’ performance on a diverse set of object-oriented programming tasks. The dataset comprises a curated collection of 100 tasks. These tasks cover a wide range of concepts, including inheritance, polymorphism, encapsulation, etc. Each coding task is in the format of the class skeleton, outlining target method description inside the class. The complexity of this dataset resides in its abstraction and hierarchical class structure. Tested models must generate large-scale code units and establish connections between each target method within the entire class, rather than focusing solely on individual functions.
The dataset provides two prompt designs for LLMs with or without IF ability. In our experiments, we employ the class skeleton as the prompt for GPT-based models, a system prompt along with task instructions for the WizardCoder.
4.1.2 Evaluation
We employ the same evaluation methodology as the original datasets for both DS-1000 and ClassEval.
  • DS-1000. we follow the original dataset using Pass@1 accuracy. This evaluation was conducted across total and four types of perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
  • ClassEval. we follow the original dataset and evaluate model performance through Pass@$K$ metric. Specifically, we calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
4.1.3 Base LLMs
This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework. GPT-4 is utilized in its gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version. To further investigate the performance of CONLINE on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al. (2023) as a base model with WizardCoder-Python-13B-V1.0 version.
4.1.4 Baselines
For the DS-1000, we selected three LLM-based frameworks as baselines.
1. DocPrompt Zhou et al. (2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools. The model then conditions on these documents, along with the problem description, to generate code.
2. Self-Debugging Chen et al. (2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors. For the sake of a fair comparison, we utilize its "simple" variant.
3. SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
For the ClassEval, We select four LLM-based code generation models as baselines: Instruct-CodeGen [^2] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al. (2023), Instruct-StarCoder [^3] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al. (2023)
It is important to note that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing. However, within the context of the real-world scenario of CONLINE, test cases from the dataset should not be used within the framework. Without these test cases, they are entirely incapable of functioning, hence we have positioned our CONLINE at a disadvantage for comparison purposes. Despite this, we are still able to achieve SOTA performance.
4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CONLINE surpasses the current SOTA model, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CONLINE exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CONLINE in generating complex code. CONLINE employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refines the details to further enhance the correctness of the code.
Table 2: Table of Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
Method Class-level Method-level
Pass@1 Pass@3 Pass@5 Pass@1 Pass@3 Pass@5
Instruct-StarCoder 10.2 12.7 14.0 23,1 26.5 27.7
SantaCoder 8.6 9.9 10.0 27.7 33.0 34.9
Instruct-CodGen 8.2 12.3 13.0 24.9 34.3 37.1
WizardCoder 12.2 20.0 23.0 35.2 47.1 51.1
CONLINE (GPT-4) 46.3 49.5 52.8 67.9 72.5 77.6
   w/o refinement of output 43.5 46.8 51.4 66.4 69.0 73.4
   w/o refinement of error 46.2 49.5 51.7 67.9 72.5 77.2
   w/o generation of test case 42.7 47.9 50.6 65.9 70.8 72.4
   w/o online retrieval 37.2 42.5 44.9 60.4 65.7 69.8
   GPT-4 36.2 39.3 43.5 58.6 64.9 67.3
CONLINE (GPT-3.5) 33.9 43.6 45.8 58.4 62.0 64.7
   w/o refinement of output 31.2 40.7 43.5 56.8 60.3 64.2
   w/o refinement of error 30.7 42.5 41.4 53.6 62.0 61.3
   w/o generation of test case 31.0 40.2 41.9 55.4 59.1 61.7
   w/o online retrieval 31.4 40.6 42.8 55.7 59.4 62.3
   GPT-3.5 27.0 32.1 35.4 52.3 54.6 59.4
For the ClassEval dataset, the results are shown in Table 2. Our experiments demonstrate that CONLINE has an overall higher porformance on both class-level and method-level Pass@$K$ evaluation. Specifically, COLINE(GPT-4) outperforms the WizardCoder (best baseline model) significantly by an average of 31.3% on Class-level and 28.3% in Method level; COLINE(GPT-3.5) outporforms the WizardCoder by an average of 22.7& in Class-level and 17.2% in Method-level.
4.3 Ablation Study
In this work, to validate the effectiveness of CONLINE, we conduct different ablation studies, with results presented in Table 1 and Table 2. Details on the ablation study are shown in §C.
CONLINE significantly enhances the base model’s ability to generate complex code. Compared to the base model, CONLINE (GPT-4) has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CONLINE.
Online search, generation of test cases, and serialization each contribute to the model’s performance improvements. Compared to CONLINE, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset. The online search improves the model by providing common code patterns, which reduces the difficulty of the model in generating initial code. Serialization, by converting inputs and outputs into a sequential format, allows the model to more intuitively observe inputs and outputs that are too lengthy or are in non-textual modalities, thereby strengthening its ability to solve complex code problems.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CONLINE shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval datset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
4.4 Analysis of Different Base Models Performance
Table 3: Table of Different Base Models Results for DS-1000. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
Method Perturbation Total/Avg.
Origin Surface Semantic Diff-Rewrite
GPT-4 64.47 69.74 56.96 43.83 60.20
   w/ retrieve 66.23 71.05 59.57 45.68 62.10
   w/ refine 68.64 70.39 60.00 51.23 64.10
   CONLINE 71.71 74.34 66.96 53.09 68.00
GPT-3.5 57.02 43.42 40.00 32.72 47.10
   w/ retrieve 47.15 25.00 36.96 25.31 37.90
   w/ refine 55.70 50.66 44.35 35.80 49.10
WizardCoder 41.01 21.71 31.74 16.05 31.90
   w/ retrieve 15.79 9.21 12.17 9.88 13.00
   w/ refine 39.69 21.71 30.00 15.43 30.80
Table 3 shows the performance results of CONLINE on the DS-1000 dataset with different base models. We can see that GPT-4 has been comprehensively improved with CONLINE, but the performance on GPT-3.5 and WizardCoder is mixed. This indicates that CONLINE requires the model to have the following capabilities to enhance its performance:
For code generation planning ability, the higher the complexity of the code that needs to be generated, the higher the demand for planning ability. Planning capability is key to online retrieval; only correct planning can generate appropriate queries to retrieve useful information. After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval as shown in Table 2, yet it decreased by 9.2% on DS-1000 as shown in Table 3. For ClassEval, we plan for each function during online retrieval. The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as that of DS-1000. Therefore, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
Code generation also necessitates the models to possess the capability of in-context learning. The generated code should be built on all the above provided contents, and the understanding of the generated results based on the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder on DS-1000 is comparatively interior to GPT models. One possible reason for this lies in the limited in-context learning capability of WizardCoder given some of prompts of DS-1000 are lengthy and complicated, which makes it difficult for WizardCoder to comprehend the related context and make corresponding code modifications accurately.
4.5 Case Study
Figure 3: Case Study for Correctness Testing.
For the case study on online retrieval, refer to Figure 4. It can be observed that by imitating the usage of functions found through online search, the model is better equipped to prepare the required parameters for the functions and to generate corresponding code. This significantly reduces the difficulty of generating complex code.
For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thereby improving subtle errors and generating the correct code.
5 Conclusion
In conclusion, CONLINE introduces a novel framework for generating complex code in real-word scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CONLINE outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
Limitations
The main limitation of our research is that it has underlying issues of exceeding the allowed times of accesses due to multiple calls to the Google Search API. Similarly, we also have made multiple API calls to test and enhance the performance of the GPT models.
Ethics Policy
This research does not pose any ethical concerns. The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
Appendix A Details for on-line search
Table 4: Website Base Station Distributions Table During the DS-1000 Online Retrieval Process.
Website Proportion
https://stackoverflow.com 57.92%
https://numpy.org 8.59%
https://pandas.pydata.org 5.70%
https://www.geeksforgeeks.org 5.07%
https://docs.scipy.org 4.76%
https://matplotlib.org 3.20%
https://www.tensorflow.org 3.04%
http://scikit-learn.org 2.42%
The Others 9.29%
Appendix B Details for Base Models
The parameter details for each model in the experiment are as follows:
  • GPT-4: model: gpt-4-32k-0613, temperature: 0, top_p: 0.95, max_tokens: 1024.
  • GPT-3.5: model: gpt-35-turbo-16k-0613, temperature: 0, top_p: 0.95, max_tokens: 1024.
  • WizardCoder: WizardCoder-Python-13B-V1.0, temperature: 0, top_p: 0.95, max_tokens: 1024.
Appendix C Details for Ablation Study
  • Without refinement of output: During the refinement process, the output result is not refined; that is, refinement is conducted solely based on the error.
  • Without refinement of error: During the refinement process, the error is not refined; that is, refinement is conducted solely based on the output result.
  • Without serialization: During the refinement process, the input and output are not serialized; instead, their printout results are directly used as input.
  • Without generation of test cases: Test cases are not generated. Since refinement cannot be performed without test cases, only online retrieval is conducted.
  • Without online retrieval: Online retrieval is not performed, and the process is limited to refinement with correctness testing.
Appendix D Case Study
Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset. Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices. After unsuccessful attempts, the model turns to an online search through a self-generated query and integrating the research results as part of next prompt. Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class. Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
Figure 4: Case Study for Online Retrieval.
Appendix E Prompts of CONLINE
Figure 5: Plan and Queries Generation Prompt on DS-1000.
Figure 6: Plan and Queries Generation Prompt on ClassEval.
Figure 7: Online Retrieval Code Generation Prompt on DS-1000 and ClassEval.
Figure 8: Generation of Test Case Prompt on DS-1000.
Figure 9: Generation of Test Case Prompt on ClassEval.
Figure 10: Refinement with Correctness Testing Code Generation Prompt on DS-1000 and ClassEval.

### v2

@@
-CONLINE: Complex Code Generation and Refinement with Online Searching and Correctness Testing
+CoCoST: Automatic Complex Code Generation with Online Searching and Correctness Testing
 Abstract
-Large Language Models (LLMs) have revolutionized code generation ability by converting natural language descriptions into executable code.
+Large Language Models have revolutionized code generation ability by converting natural language descriptions into executable code.
 However, generating complex code within real-world scenarios remains challenging due to intricate structures, subtle bugs, understanding of advanced data types, and lack of supplementary contents.
-To address these challenges, we introduce the CONLINE framework, which enhances code generation by incorporating planned online searches for information retrieval and automated correctness testing for iterative refinement.
-CONLINE also serializes the complex inputs and outputs to improve comprehension and generate test case to ensure the framework’s adaptability for real-world applications.
-CONLINE is validated through rigorous experiments on the DS-1000 and ClassEval datasets.
-It shows that CONLINE substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality and reliability of LLMs in generating intricate code.
-CONLINE: Complex Code Generation and Refinement
+To address these challenges, we introduce the CoCoST framework, which enhances complex code generation by online searching for more information with planned queries and correctness testing for code refinement.
+Moreover, CoCoST serializes the complex inputs and outputs to improve comprehension and generates test cases to ensure the adaptability for real-world applications.
+CoCoST is validated through rigorous experiments on the DS-1000 and ClassEval datasets.
+Experimental results show that CoCoST substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality of LLMs in generating complex code.
+CoCoST: Automatic Complex Code Generation
 with Online Searching and Correctness Testing
@@
 1 Introduction
-Figure 1: An Example of the Human Programmer Code-writing Process Imitated by the CONLINE.
+Figure 1: An Example of the Human Developer Code-writing Process Imitated by the CoCoST.
 After the problem is received, an online search is performed to simulate search results and create an initial version of code.
@@
 The code is refined based on the correctness of these results.
-Automated code generation from natural language descriptions has became more realistic, as large language models (LLMs) show their potential to generate complex and accurate code Manna and Waldinger (1971); Li et al.
+Automatic code generation from natural language descriptions is becoming more realistic, as large language models (LLMs) show their potential to generate accurate code Li et al.
 (2023); Luo et al.
@@
 (2024).
-There are a variety of methods that have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al.
+Various methods have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al.
 (2023); Jiang et al.
@@
 (2023).
-In real-world scenarios, complex code generation can involve intricate problem description, sophisticated code logic, and advanced data types Lai et al.
+However, complex code generation is a more difficult task, which involves intricate problem description, sophisticated code logic, and advanced data types Lai et al.
 (2022); Du et al.
@@
 (2023).
-For example (see Figure 1), solving the required code generation involves multiple transformations of various matrices, as well as the invocation of functions that demands numerous and stringent parameters.
-Existing methods were not enough to address the following new challenges that emerge during complex code generation:
-  1. 1.
-    Complex code generation is hindered by intricate structures and logic. E.g., the complexity of the example in Figure 1 is as previously described.
-  2. 2.
-    Obscure bugs in complex code necessitate critical detection and refinement.
-Existing methods often refine code based on execution errors testing Zhang et al.
+The existing methods struggle to address the arising challenges:
+Challenge 1: Offline documents cannot meet the diverse demands of code generation.
+In real-world scenarios, these demands often exceed the capabilities of limited offline documents.
+For example, problem descriptions may involve functions that are not covered by pre-collected documents.
+Additionally, complex code generation for diverse needs often entails highly complex logic and a series of transformation functions like the programming problem in Figure 1, where simple API examples in documents fail to provide adequate guidance.
+Challenge 2: In real-world situations, there is often a shortage of test cases (e.g., test cases in Figure 1) for automatic code generation.
+Most existing work depends heavily on pre-existing test cases in datasets Zhang et al.
 (2023); Jiang et al.
-(2023), but for complex codes, sometimes the executable code has non-obvious bugs,e.g., initial code in Figure 1.
-  3. 3.
-    The complexity of inputs and outputs hampers LLM’s comprehension ability during code refinement.
-In complex code, inputs and outputs like Pandas DataFrames, PyTorch tensors, and matplotlib PNG images are intricate.
-Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities and so on.
-  4. 4.
-    Real-world scenarios frequently lack supplemental content, such as test cases and documentation pools.
-Existing works predominantly rely on pre-existing test cases in datasets Zhang et al.
+(2023) , which are difficult to acquire directly in practical scenarios.
+Challenge 3: Hidden bugs in complex code require meticulous identification and refinement.
+Current techniques frequently enhance code by analyzing execution errors Zhang et al.
 (2023); Jiang et al.
-(2023) or off-line document pools Zhou et al.
-(2023); Jiang et al.
-(2023), which are challenging to obtain directly in real-world scenarios.
-To address the above challenges, we propose a new code generation framework CONLINE that enhances the code generation and refinement quality of LLMs with planned online searching and automated correctness testing steps in this paper.
-The intuition of CONLINE is straightforward: During the process of coding, most human programmers are not bothered by the above challenges as shown in Figure 1.
-They could easily bypass them by searching on-line for answers, experiences, or guidelines.
-Also, they are able to design test cases and run code against them to verify the correctness of the code logic.
-Correspondingly, CONLINE uses online search for code generation to partially address Challenges 1 and 4.
-For Challenge 1, it offers following advantages: (1) Retrieving information from library documents allows the retrieval of standardized function syntax, preventing errors from model hallucinations e.g., generating non-existing parameters.
-(2) Retrieving information from the latest blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code.
-For Challenge 4, online search encompasses the content of offline document pools and extends beyond with its superior capability to adapt to the diverse and evolving needs of real-world scenarios.
-Directly searching for solutions to such problems is difficult, because the problems are intricate and may involve multiple challenges.
-Therefore, we propose online search with generation query through planning.
-To address the Challenge 2, this work proposes correctness testing in refinement.
-LLMs may generate complex code that contains subtle errors, resulting in code that is executable but does not produce the desired output.
-Therefore, during the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors.
-We propose incorporating both the execution output results and the errors as integral parts of the refinement to enhance the correctness of the code.
-In refinement with correctness testing, we propose generation of test cases to address the Challenge 3.
-Test cases are crucial, as they serve as indispensable inputs for the code execution.
-This work utilizes LLMs to automatically generate test cases, making the framework applicable even in real-world scenarios where no known test cases exist.
-To address the Challenge 4, we propose serialization of input and output.
-It renders the input and output more intuitive and understandable for the model, particularly for those that are excessively long or non-textual modalities.
-We test the effectiveness of CONLINE on two complex code generation datasets using three different LLMs.
-For the complex code generation datasets, we selected DS-1000, which focuses on complex data science code generation, and the Class Eval dataset, which is oriented towards class-level code generation.
-By comparing with the existing state-of-the-art (SOTA) baseline, we achieved a 10.9% improvement on DS-1000 and average 9.47% on ClassEval, confirming the effectiveness of CONLINE.
-Through comprehensive ablation studies, we verified the effectiveness of each component of CONLINE.
-Moreover, given the varying performance of CONLINE on different base models, we discovered that CONLINE requires models to have capabilities such as Planning, which vary according to the complexity of the problem, in order to further enhance the model’s code generation abilities.
+(2023).
+But in the case of complex code, the executable code sometimes contains hidden bugs like the highlighted part of the initial code in Figure 1.
+To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be publicly available upon acceptance.
+(Automatic Complex Code Generation with Online Searching and Correctness Testing) that improves the generation and refinement of complex code by LLMs through the planned online searching and automatic correctness testing steps.
+The intuition of CoCoST is straightforward: During the coding process, most human developers are not bothered by the above challenges, as illustrated in Figure 1.
+Developers can easily overcome these obstacles by searching online through engines (e.g., Google and Bing) for solutions, experiences, and guidelines.
+In addition, they can create test cases and execute code to ensure the correctness of the code logic.
+To address Challenges 1, CoCoST proposes an online search methodology.
+This process involves querying web search engines and then extracting pertinent information to construct prompts for LLMs.
+The approach presents several benefits: (1) Retrieving information from the up-to-date blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code.
+(2) Online search extends beyond the scope of static offline documentation, covering a wider array of problems without being confined to a predetermined set.
+Meanwhile, it reduces the effort developers need to expend in assembling documentation, thereby increasing the framework’s level of automation.
+However, using problem descriptions as search queries can be challenging due to the complexity of the problems, which often encompass multiple issues.
+Therefore, we propose an online search with generation query through planning.
+To address the Challenge 2, we introduce generation of test cases during refinement.
+Several studies Chen et al.
+(2022); Shinn et al.
+(2023) have attempted to generate tests.
+However, these methods often fall short when applied to the generation of complex code due to its intricate logic and outputs, which complicate the direct production of accurate tests (both the inputs and expected outputs for the solution code).
+CoCoST utilizes LLMs to automatically generate test cases (the inputs for the code).
+This strategy cleverly focuses on generating test cases without attempting to produce complete tests.
+It significantly simplifies the process of test case generation and facilitates its precise creation for complex code.
+To address the Challenge 3, this work prioritizes correctness testing in refinement.
+During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors.
+CoCoST incorporate both the execution output results and the errors into within the refinement prompts for LLMs to enhance the correctness.
+Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, Matplotlib charts.
+Thus, CoCoST propose serialization of input and output to convert them into understandable sequences before being processed by LLMs.
+Particularly those are excessively long or non-textual modalities.
+We evaluate the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval).
+By comparing with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and average 9.47% on ClassEval.Moreover, we analyze and discovery that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem.
+In summary, our main contributions are:
+  • We propose the novel CoCoST framework to generate complex code, and can be automatic in real-world scenarios.
+  • We apply online search in code generation for the first time to our knowledge, in order to generate the complex code.
+  • We prioritize correctness testing in refinement with test case generation and serialization of input and output, in order to refine hidden bugs in complex code.
+  • We conduct experiments on DS-1000 and ClassEval datasets anddemonstrate the effectiveness and universality of CoCoST.
 2 Related Work
+Code generation datasets.
 The realm of automated code generation has been propelled by benchmark datasets such as HumanEval Chen et al.
@@
 These endeavors reflect an emerging trend towards enhancing models’ abilities to produce sophisticated and domain-specific code structures.
-Recent research has concentrated on improving Large Language Models (LLMs) for code generation via retrieve, refine and so on strategies.
-Retrieve enhances model performance by pulling in external information, with techniques like DocPrompt Zhou et al.
-(2023) accessing fixed document repositories and SELVEVOLVE Jiang et al.
-(2023) using models as knowledge bases.
-However, these are limited to document-sourced content and provide limited benefits.
+In this paper, we select datasets with complex code generation to evaluate CoCoST.
+Retrieval-augmented code generation.
+With the emergence of Large Language Models (LLMs), a variety of retrieval-augmented techniques have been developed to compensate for issues such as the inherent knowledge limitations.
+DocPromptZhou et al.
+(2023) and SELVEVOLVEJiang et al.
+(2023) leverage document libraries or models as knowledge bases to improve code generation.
+However, their reliance on fixed document libraries limits the scope of information they can provide and confines the generated code to the context of these libraries.
+Furthermore, the prerequisite of preestablished document libraries prevents these approaches from being fully autonomous in real-world frameworks.
+Solutions such as WebGPT Nakano et al.
+(2022), LaMDA Thoppilan et al.
+(2022), and FreshLLMs Vu et al.
+(2023) enhance the performance of natural language tasks by using online search or open web knowledge.
+However, because complex code generation often involves multiple steps and complexities, these methods struggle with direct application to complex code generation.
+Code refinement.
 Refine iteratively enhances generated code for greater precision.
-Projects such as Self-Debug Chen et al.
-(2023) and SELFEVOLVE Jiang et al.
-(2023) focus on error refinement, while Self-EditZhang et al.
-(2023) uses code execution results as feedback.
-Despite advancements, these methods often miss subtle bugs and are less applicable to complex, real-world code generation without readily available test cases or ground truths.
-Figure 2: The Pipline of CONLINE.
+Self-DebugChen et al.
+(2023), SELFEVOLVEJiang et al.
+(2023), and Self-EditZhang et al.
+(2023) improve code generation by refining code through the resolution of errors identified during execution.
+These methods effectively address errors, while when it comes to complex code generation, subtle bugs also play a significant role in the overall error landscape.
+Moreover, relying on pre-existing tests from datasets in refinement limits their autonomy in real-world applications, where such tests may not be readily available.
+CodeT Chen et al.
+(2022), Reflexion Shinn et al.
+(2023), and CODECHAIN Le et al.
+(2023) seek to strengthen code generation by creating tests.
+But the tests they generate include not only the inputs for the solution code but also the expected outputs.
+This poses a substantial challenge for complex code generation, where the logic can be intricate and certain problems may not lend to straightforward ground truth generation.
+Figure 2: The Pipline of CoCoST.
 Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps.
@@
 In this work, we adopt a two-step approach for code generation, mirroring the way humans write code.
-The first step is retrieve, where relevant information is obtained through online saerch and utilized by LLMs to generate initial code.
-The second step is refine, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
-3.1 Retrieve
+The first step is retrieval, where relevant information is obtained through online search and utilized by LLMs to generate initial code.
+The second step is refinement, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
+3.1 Retrieval
 The difficulty in achieving effective online retrieval lies in formulating optimal search queries.
@@
 Directly searching for solutions to such problems is inaccurate and difficult.
-On the other hand, it is challenging that matching queries directly through methods for offline document like similarity calculations, due to the nature of online libraries.
-Therefore, we propose using planning to solve the challenge.
+On the other hand, it is challenging that match queries directly through methods for offline documents like similarity calculations, due to the nature of online libraries.
+So we propose generating queries through planning to solve the challenge.
 The retrieval process is divided into three steps: 1.
@@
 3.
-The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $I$:
-$\widehat{W}_{0}\sim p_{\theta}(.|D,I)$ (1)
+The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $INFO$:
+$\widehat{W}_{0}\sim p_{\theta}(.|D,INFO)$ (1)
 3.1.1 Generation Query through Planning
@@
 3.1.2 Online Search
-For the above generated queries, we conduct an online search. We use the online search API [^1] https://github.com/Nv7-GitHub/googlesearch for the search process.
-$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$
-where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ and $N_{u}=1$.
+For the above generated queries, we conduct an online search.
+In this study, we use the online search API [^2] https://github.com/Nv7-GitHub/googlesearch for the search process as Equation (3).
+CoCoST can also be applied to private or domain-specific knowledge repositories as long as they are accessible via query, with details in §A.
+$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$ (3)
+where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ , $N_{u}=1$.
 Through the analysis of the website distribution Table 4, we observed that more than 90% of the URLs are concentrated on a total of 8 websites.
@@
 The information $INFO$ is composed of details from each query $q_{j}$, each url $url_{k}$, with each piece of information $info_{j,k}$ extracted.
-3.2 Refine
+3.2 Refinement
 Existing work Chen et al.
 (2023); Jiang et al.
-(2023) typically emphasizes correction of errors identified during the refinement process.
+(2023) typically emphasizes the correctness of errors identified during the refinement process.
 However, we observe that refining code that produces error-free outputs is equally crucial during the refinement process.
-Therefore, we introduce correction testing in §3.2.1.
+Therefore, we introduce correctness testing in §3.2.1.
 Additionally, we propose methods for the generation of test cases and serialization of inputs and outputs during the refinement process.
@@
 Incorporating both the error and the output result into the refinement process allows the model to take advantage of self-correction mechanisms.
-$\left\{\begin{array}[]{ccl}e_{j,k},o_{j,k}&=&execute(W_{j},i_{k}),j\in\{1,\ldots,N_{f}\}\\ INFO_{e_{j,k}}&=&\{e_{j,k},extract(search(e_{j,k}))\}\\ \widehat{W}_{j+1}&\sim&p_{\theta}(.|D,W_{j},\{S_{i},S_{o_{j}},INFO_{e_{j}}\}_{k}),\\ &&\hfil k\in\{1,\ldots,N_{i}\}\end{array}\right.$
+[…]
 where, $N_{f}$ is the total number of refinement steps, $N_{i}$ is the number of inputs.
-$i_{K}$ is the k-th input for the problem from Equation (3), $S_{i}$ and $S_{o_{j}}$ is the serialization of the input and output from Equation (4).
-3.2.2 Generation of Test Case
-Test cases is crucial, as they serve as indispensable inputs for the code execution in refinement.
+$i_{K}$ is the k-th input for the problem from Equation (4), $S_{i}$ and $S_{o_{j}}$ is the serialization of input and output from Equation (5).
+3.2.2 Generation of Test Cases
+Test cases are crucial, as they serve as indispensable inputs for the code execution in refinement.
 While, existing works in refining code predominantly rely on pre-existing test cases in datasets Zhang et al.
@@
 (2023) even uses the ground truth output of the test case to refine the code, which is even more challenging to obtain for complex code problems in real-world scenarios.
-Because their problems involve various logical operations, and even with input test cases, deriving answers directly without code-based computations is demanding.
-In this study, we propose generation of test case with LLMs to adapt to real-world scenarios.
+Because their problems involve various logical operations, deriving answers directly without code-based computations is demanding.
+CoCoST introduces generation of test cases with LLMs to adapt to real-world scenarios.
 $\left\{\begin{array}[]{ccl}\widehat{I}&\sim&p_{\theta}(.|D)\\ I&=&\{i_{1},\ldots,i_{N_{i}}\}\end{array}\right.$ (4)
@@
   1. 1.
-    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes a truncated string of data, data type, data shape, and statistical information.
+    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes data truncated string, data type, data shape, and statistical information.
   2. 2.
@@
 $S_{n}=serialize(n),n\in\{i_{k},o_{j,k}\}$ (5)
-Table 1: Table of Main Results and Ablation Study for DS-1000.
-The base model for CONLINE is GPT-4.
-All metric numbers are represented as percentages.
+Table 1: Main Results and Ablation Study for DS-1000.
+The base model for CoCoST is GPT-4.
+All metrics are represented as percentages.
 For each metric, the bold number indicates the highest performance.
@@
 Origin Surface Semantic Diff-Rewrite
+Codex 44.93 37.94 34.35 16.94 39.20
 DocPrompting 53.95 50.00 38.39 21.05 43.30
@@
 SELFEVOLVE 66.23 67.11 48.70 33.95 57.10
-CONLINE 71.71 74.34 66.96 53.09 68.00
+Reflexion 58.99 73.03 52.17 48.77 57.90
+CoCoST 71.71 74.34 66.96 53.09 68.00
      w/o refinement of output 68.42 69.74 62.61 48.77 64.10
@@
      w/o online retrieval 68.64 70.39 60.00 51.23 64.10
-   GPT-4 64.47 69.74 56.96 43.83 60.20
+     w/o all (GPT-4 only) 64.47 69.74 56.96 43.83 60.20
 4 Experiment
@@
 4.1.1 Datasets
-We conduct experience on two complex code generation datasets:
-1.
+We conduct experiences on two complex code generation datasets:
 DS-1000 Lai et al.
@@
 Secondly, the input-output involves various complex data structures related to data science, making the code logic intricate during transformations of the data.
-The dataset provides both Insertion and Completion style prompts, where the data is the same, differing only in prompt format, thus yielding similar results.
-In this paper, experiments are conducted with the Completion style prompt.
-2.
+Further details of DS-1000 implementation are shown in §B.1.
 ClassEval Du et al.
@@
 These tasks cover a wide range of concepts, including inheritance, polymorphism, encapsulation, etc.
-Each coding task is in the format of the class skeleton, outlining target method description inside the class.
+Each coding task is in the format of the class skeleton, outlining the target method description inside the class.
 The complexity of this dataset resides in its abstraction and hierarchical class structure.
@@
 In our experiments, we employ the class skeleton as the prompt for GPT-based models, a system prompt along with task instructions for the WizardCoder.
-4.1.2 Evaluation
-We employ the same evaluation methodology as the original datasets for both DS-1000 and ClassEval.
-  • DS-1000. we follow the original dataset using Pass@1 accuracy. This evaluation was conducted across total and four types of perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
-  • ClassEval. we follow the original dataset and evaluate model performance through Pass@$K$ metric. Specifically, we calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
-4.1.3 Base LLMs
-This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework.
-GPT-4 is utilized in its gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version.
-To further investigate the performance of CONLINE on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al.
-(2023) as a base model with WizardCoder-Python-13B-V1.0 version.
-4.1.4 Baselines
-For the DS-1000, we selected three LLM-based frameworks as baselines.
-1.
-DocPrompt Zhou et al.
-(2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools.
-The model then conditions on these documents, along with the problem description, to generate code.
-2.
-Self-Debugging Chen et al.
-(2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors.
-For the sake of a fair comparison, we utilize its "simple" variant.
-3. SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
-For the ClassEval, We select four LLM-based code generation models as baselines: Instruct-CodeGen [^2] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al.
-(2023), Instruct-StarCoder [^3] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al.
-(2023)
-It is important to note that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing.
-However, within the context of the real-world scenario of CONLINE, test cases from the dataset should not be used within the framework.
-Without these test cases, they are entirely incapable of functioning, hence we have positioned our CONLINE at a disadvantage for comparison purposes.
-Despite this, we are still able to achieve SOTA performance.
-4.2 Main Results
-Regarding the DS-1000 dataset, the main results are shown in Table 1.
-CONLINE surpasses the current SOTA model, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA.
-Especially under the Diff-Rewrite perturbation setting, CONLINE exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CONLINE in generating complex code.
-CONLINE employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refines the details to further enhance the correctness of the code.
-Table 2: Table of Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
+Table 2: Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
 Method Class-level Method-level
 Pass@1 Pass@3 Pass@5 Pass@1 Pass@3 Pass@5
-Instruct-StarCoder 10.2 12.7 14.0 23,1 26.5 27.7
+Instruct-StarCoder 10.2 12.7 14.0 23.1 26.5 27.7
 SantaCoder 8.6 9.9 10.0 27.7 33.0 34.9
@@
 WizardCoder 12.2 20.0 23.0 35.2 47.1 51.1
-CONLINE (GPT-4) 46.3 49.5 52.8 67.9 72.5 77.6
+Reflexion 24.1 30.7 35.2 43.4 51.6 61.8
+CoCoST 46.3 49.5 52.8 67.9 72.5 77.6
      w/o refinement of output 43.5 46.8 51.4 66.4 69.0 73.4
@@
      w/o online retrieval 37.2 42.5 44.9 60.4 65.7 69.8
-   GPT-4 36.2 39.3 43.5 58.6 64.9 67.3
-CONLINE (GPT-3.5) 33.9 43.6 45.8 58.4 62.0 64.7
-   w/o refinement of output 31.2 40.7 43.5 56.8 60.3 64.2
-   w/o refinement of error 30.7 42.5 41.4 53.6 62.0 61.3
-   w/o generation of test case 31.0 40.2 41.9 55.4 59.1 61.7
-   w/o online retrieval 31.4 40.6 42.8 55.7 59.4 62.3
-   GPT-3.5 27.0 32.1 35.4 52.3 54.6 59.4
+     w/o all (GPT-4 only) 36.2 39.3 43.5 58.6 64.9 67.3
+4.1.2 Evaluation
+We employ the same evaluation methodology as the original datasets for both DS-1000 and ClassEval.
+DS-1000. We follow the original dataset using Pass@1 accuracy. This evaluation is conducted across total and perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
+ClassEval. We follow the original dataset using Pass@$K$ metric. We calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
+4.1.3 Base LLMs
+This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework.
+GPT-4 is utilized in gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version.
+To further investigate the performance of CoCoST on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al.
+(2023) as a base model with WizardCoder-Python-13B-V1.0 version.
+4.1.4 Baselines
+For the DS-1000, we selected four LLM-based frameworks as baselines: DocPrompt Zhou et al.
+(2023), Self-Debugging Chen et al.
+(2023), SELFEVOLVE Jiang et al.
+(2023) and Reflexion Shinn et al.
+(2023).
+DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools.
+Self-Debugging depends on a Python interpreter to instruct language models in revising Python code containing errors.
+SELFEVOLVE employs LLMs as both sources of knowledge and self-reflective programmers.
+Reflexion utilizes reflective feedback with generated tests and episodic memory to process task feedback.
+Details are shown in §B.3.
+For the ClassEval, we select five LLM-based code generation models and frameworks as baselines: Instruct-CodeGen [^3] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al.
+(2023), Instruct-StarCoder [^4] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al.
+(2023) and Reflexion Shinn et al.
+(2023).
+4.2 Main Results
+Regarding the DS-1000 dataset, the main results are shown in Table 1.
+CoCoST surpasses the current SOTA framework, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA.
+Especially under the Diff-Rewrite perturbation setting, CoCoST exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CoCoST in generating complex code.
+CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refining the details to further enhance the correctness of the code.
 For the ClassEval dataset, the results are shown in Table 2.
-Our experiments demonstrate that CONLINE has an overall higher porformance on both class-level and method-level Pass@$K$ evaluation.
-Specifically, COLINE(GPT-4) outperforms the WizardCoder (best baseline model) significantly by an average of 31.3% on Class-level and 28.3% in Method level; COLINE(GPT-3.5) outporforms the WizardCoder by an average of 22.7& in Class-level and 17.2% in Method-level.
+Our experiments demonstrate that CoCoST has an overall higher performance on both class-level and method-level Pass@$K$ evaluation.
+Specifically, CoCoST outperforms the Reflexion (best baseline model) significantly by an average of 19.5% and 20.4% on the Class and Method level.
 4.3 Ablation Study
-In this work, to validate the effectiveness of CONLINE, we conduct different ablation studies, with results presented in Table 1 and Table 2. Details on the ablation study are shown in §C.
-CONLINE significantly enhances the base model’s ability to generate complex code.
-Compared to the base model, CONLINE (GPT-4) has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CONLINE.
+Table 3: Different Base Models Results for DS-1000 and ClassEval. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
+Method DS-1000 ClassEval
+Origin Surface Semantic Diff-Rewrite Total/Avg. Class-level Method-level
+GPT-4 64.47 69.74 56.96 43.83 60.20 43.5 67.3
+     + retrieve 66.23 71.05 59.57 45.68 62.10 50.6 72.4
+     + refine 68.64 70.39 60.00 51.23 64.10 44.9 69.8
+     CoCoST 71.71 74.34 66.96 53.09 68.00 52.8 77.6
+GPT-3.5 57.02 43.42 40.00 32.72 47.10 35.4 59.4
+     + retrieve 47.15 25.00 36.96 25.31 37.90 41.9 61.7
+     + refine 55.70 50.66 44.35 35.80 49.10 42.8 62.3
+     CoCoST - - - - - 45.8 64.7
+WizardCoder 41.01 21.71 31.74 16.05 31.90 23.0 51.1
+     + retrieve 15.79 9.21 12.17 9.88 13.00 18.2 41.8
+     + refine 39.69 21.71 30.00 15.43 30.80 22.3 50.7
+In this work, to validate the effectiveness of CoCoST, we conduct different ablation studies, with results presented in Table 1 and 2. Details on the ablation study are shown in §B.4.
+CoCoST significantly enhances the base model’s ability to generate complex code.
+Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
 Online search, generation of test cases, and serialization each contribute to the model’s performance improvements.
-Compared to CONLINE, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset.
+Compared to CoCoST, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset.
 The online search improves the model by providing common code patterns, which reduces the difficulty of the model in generating initial code.
 Serialization, by converting inputs and outputs into a sequential format, allows the model to more intuitively observe inputs and outputs that are too lengthy or are in non-textual modalities, thereby strengthening its ability to solve complex code problems.
+Online search outperforms offline retrieval in effectiveness and has a wider range of applicability.
+As shown in Table 1, using only online retrieval (the row w/o generation of test case) outperforms DocPrompting, which is an offline retrieval approach.
+Moreover, in real-world scenarios, as opposed to specific datasets, the types of problems encountered are more diverse.
+The scalability of online retrieval enables them to effectively address a wide range of problems.
+However, offline retrieval systems struggle to encompass all relevant information comprehensively.
 During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important.
-After separately conducting ablation studies on the output result and error, CONLINE shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval datset.
+After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval dataset.
 This indicates that the output result contributes more to the refinement process than the error.
@@
 4.4 Analysis of Different Base Models Performance
-Table 3: Table of Different Base Models Results for DS-1000. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
-Method Perturbation Total/Avg.
-Origin Surface Semantic Diff-Rewrite
-GPT-4 64.47 69.74 56.96 43.83 60.20
-   w/ retrieve 66.23 71.05 59.57 45.68 62.10
-   w/ refine 68.64 70.39 60.00 51.23 64.10
-   CONLINE 71.71 74.34 66.96 53.09 68.00
-GPT-3.5 57.02 43.42 40.00 32.72 47.10
-   w/ retrieve 47.15 25.00 36.96 25.31 37.90
-   w/ refine 55.70 50.66 44.35 35.80 49.10
-WizardCoder 41.01 21.71 31.74 16.05 31.90
-   w/ retrieve 15.79 9.21 12.17 9.88 13.00
-   w/ refine 39.69 21.71 30.00 15.43 30.80
-Table 3 shows the performance results of CONLINE on the DS-1000 dataset with different base models.
-We can see that GPT-4 has been comprehensively improved with CONLINE, but the performance on GPT-3.5 and WizardCoder is mixed.
-This indicates that CONLINE requires the model to have the following capabilities to enhance its performance:
+Table 3 shows the performance results of CoCoST on the DS-1000 dataset with different base models.
+We can see that GPT-4 has been comprehensively improved with CoCoST, but the performance on GPT-3.5 and WizardCoder is mixed.
+This indicates that CoCoST requires the model to have the following capabilities to enhance its performance:
 For code generation planning ability, the higher the complexity of the code that needs to be generated, the higher the demand for planning ability.
 Planning capability is key to online retrieval; only correct planning can generate appropriate queries to retrieve useful information.
-After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval as shown in Table 2, yet it decreased by 9.2% on DS-1000 as shown in Table 3.
-For ClassEval, we plan for each function during online retrieval.
-The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as that of DS-1000.
-Therefore, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
-Code generation also necessitates the models to possess the capability of in-context learning.
-The generated code should be built on all the above provided contents, and the understanding of the generated results based on the preceding input prompt is of great importance in the refinement stage.
+After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval, yet it decreased by 9.2% on DS-1000 as shown in Table 3.
+The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as DS-1000.
+Thus, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
+Code generation necessitates models to have in-context learning abilities.
+The generated code should be built on all the above provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage.
 In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively.
-And the overall performance of WizardCoder on DS-1000 is comparatively interior to GPT models.
-One possible reason for this lies in the limited in-context learning capability of WizardCoder given some of prompts of DS-1000 are lengthy and complicated, which makes it difficult for WizardCoder to comprehend the related context and make corresponding code modifications accurately.
-4.5 Case Study
-Figure 3: Case Study for Correctness Testing.
+And the overall performance of WizardCoder is comparatively interior to GPT models.
+This could be due to WizardCoder’s limited in-context learning ability, especially with the complex and lengthy prompts, hindering accurate context comprehension and code modification.
+4.5 Case Study and Error Analysis
 For the case study on online retrieval, refer to Figure 4.
@@
 It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results.
-The model refines the code based on the output, thereby improving subtle errors and generating the correct code.
+The model refines the code based on the output, thus improving hidden errors and generating the correct code.
+Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors.
+For the former, a discussion is provided in §C.2, while for the latter, an error analysis is conducted in §C.3.
 5 Conclusion
-In conclusion, CONLINE introduces a novel framework for generating complex code in real-word scenarios by emulating human coding processes like online searching and test case creation.
+In conclusion, CoCoST introduces a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation.
 It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data.
 The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding.
-Tested on various datasets, CONLINE outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
+Tested on various datasets, CoCoST outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
 Limitations
@@
 This research does not pose any ethical concerns. The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
-Appendix A Details for on-line search
+Appendix A Online Searching Detail
+Website Base Station Distributions Table During the DS-1000 Online Retrieval Process:
 Table 4: Website Base Station Distributions Table During the DS-1000 Online Retrieval Process.
@@
 The Others 9.29%
-Appendix B Details for Base Models
+Moreover, CoCoST can be applied to specialized, proprietary, or domain-specific knowledge repositories as long as they are accessible via query.
+Moreover, implementing queries for private datasets is easily achievable and a growing trend in data management.
+Major companies such as Google and Microsoft already offer products designed to search private data; for example, Google Workspace’s Cloud Search provides powerful capabilities for enterprises to search their private data.
+In this paper, to validate the effectiveness of our framework, we conducted tests on public online searches.
+Moving forward, the framework can be applied to an even broader range of knowledge repositories.
+Appendix B Experiment
+B.1 Datasets Detail
+Further details of DS-1000 implementation are as follows:
+  • The dataset provides both Insertion and Completion style prompts, where the data is the same, differing only in prompt format, thus yielding similar results.
+In this paper, experiments are conducted with the Completion style prompt.
+  • We implement a filtering approach to prevent data leakage and model replication of existing solutions from Stack Overflow.
+The DS-1000 dataset originates from Stack Overflow, and concurrently, over 50% of the websites we encountered during our online searches are from Stack Overflow.
+Thus, to prevent data leakage, when conducting online searching, we filter out all Stack Overflow problems belonging to the source of the DS-1000 dataset by using the Stack Overflow question_id.
+B.2 Base Models
 The parameter details for each model in the experiment are as follows:
@@
   • WizardCoder: WizardCoder-Python-13B-V1.0, temperature: 0, top_p: 0.95, max_tokens: 1024.
-Appendix C Details for Ablation Study
+B.3 Baselines Details
+  • DocPrompt Zhou et al.
+(2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools.
+The model then conditions on these documents, along with the problem description, to generate code.
+  • Self-Debugging Chen et al.
+(2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors.
+For the sake of a fair comparison, we utilize its "simple" variant.
+  • SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
+  • Reflexion Shinn et al.
+(2023): Reflexion utilize reflective feedback with generated tests and episodic memory to process task feedback.
+For the sake of a fair comparison, we utilize GPT-4 as base model and set trail number = 1.
+It is worth noting that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing.
+However, within the context of the real-world scenario of CoCoST, test cases from the dataset should not be used within the framework.
+Without these test cases, they are entirely incapable of functioning.
+B.4 Ablation Study Details
   • Without refinement of output: During the refinement process, the output result is not refined; that is, refinement is conducted solely based on the error.
@@
   • Without online retrieval: Online retrieval is not performed, and the process is limited to refinement with correctness testing.
-Appendix D Case Study
+    Appendix C Experimantal Results
+    C.1 DS-1000 Results
+    The main results for different packages in DS-1000 are shown in Table 5.
+    Table 5: Table of Main Results for different packages in DS-1000. All metric numbers are represented as percentages. The bold number indicates the highest performance.
+    Method Pandas Numpy Matplotlib Tensorflow Scipy Sklearn Pytorch Total/Avg.
+    CoCoST 59.45 75.91 75.48 71.11 61.32 63.48 77.94 68.00
+         + retrieve 51.89 70.91 68.39 66.67 52.83 70.43 60.29 62.10
+         + refine 55.67 72.73 74.19 64.44 54.72 60.00 70.59 64.10
+         GPT-4 only 52.23 70.45 67.74 55.56 50.00 64.35 55.88 60.20
+    The results indicate that CoCoST shows a more pronounced effect on libraries whose inputs and outputs are more complex or more challenging for LLMs to intuitively understand, such as Matplotlib, TensorFlow, and PyTorch.
+On Sklearn, CoCoST experiences a slight decline due to its test cases containing complex objects, which present a significant challenge in generating test cases.
+Consequently, CoCoST’s performance on Sklearn is not as strong as with the other libraries.
+    C.2 Analysis of Pipeline
+    Regarding the generation of test cases, to prove that the generated test cases are comparable to the ground truth test cases, we substitute the generated test cases in CoCoST with ground truth test cases and conduct experiment on the DS-1000 dataset using GPT-4.
+The results showed that the performance was 68.70%, only marginally higher by 0.7% compared to the use of generated test cases (68.00%).
+This proves that the effects of both are comparable and that the generated test cases rarely lead to errors.
+    Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval.
+On the DS-1000 dataset using GPT-4, this approach result in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%).
+This confirms that the generated plans are significantly effective.
+    C.3 Error Analysis
+    For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements.
+Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect.
+For the former cases, it shows that CoCoST is able to fix some bad cases in the refinement stage even though the retrieval contents have some errors.
+For the later cases, we do observe some cases that are worthy of further research as potential directions for future work.
+First, the search content could be more detailed.
+E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but is exactly where the bug in the code.
+Second, the search query could be more targeted.
+The descriptions of some queries are not specific enough in terms of some complicated problems.
+Therefore, more sub-queries are needed to help model receive clearer instructions.
+C.4 Case Study
 Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset.
 Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices.
-After unsuccessful attempts, the model turns to an online search through a self-generated query and integrating the research results as part of next prompt.
+After unsuccessful attempts, the model turns to an online search through a self-generated query and integrates the research results as part of the next prompt.
 Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class.
 Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
+Figure 3: Case Study for Correctness Testing.
 Figure 4: Case Study for Online Retrieval.
-Appendix E Prompts of CONLINE
+Appendix D Prompts of CoCoST
 Figure 5: Plan and Queries Generation Prompt on DS-1000.

### v3

@@
 Figure 1: An Example of the Human Developer Code-writing Process Imitated by the CoCoST.
-After the problem is received, an online search is performed to simulate search results and create an initial version of code.
+After the problem is received, an online search is performed to simulate search results and create an initial version of the code.
 Test cases are then generated, and the code is executed to produce output results.
@@
 But in the case of complex code, the executable code sometimes contains hidden bugs like the highlighted part of the initial code in Figure 1.
-To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be publicly available upon acceptance.
+To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be open-sourced on https://github.com/microsoft/CoCoST.
 (Automatic Complex Code Generation with Online Searching and Correctness Testing) that improves the generation and refinement of complex code by LLMs through the planned online searching and automatic correctness testing steps.
@@
 In addition, they can create test cases and execute code to ensure the correctness of the code logic.
-To address Challenges 1, CoCoST proposes an online search methodology.
-This process involves querying web search engines and then extracting pertinent information to construct prompts for LLMs.
+To address Challenge 1, CoCoST proposes an online search methodology.
+This process involves querying web search engines and then extracting pertinent information to construct LLM prompts.
 The approach presents several benefits: (1) Retrieving information from the up-to-date blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code.
-(2) Online search extends beyond the scope of static offline documentation, covering a wider array of problems without being confined to a predetermined set.
+(2) Online search extends beyond the scope of static offline documentation, covering a wider range of problems without being confined to a predetermined set.
 Meanwhile, it reduces the effort developers need to expend in assembling documentation, thereby increasing the framework’s level of automation.
-However, using problem descriptions as search queries can be challenging due to the complexity of the problems, which often encompass multiple issues.
-Therefore, we propose an online search with generation query through planning.
-To address the Challenge 2, we introduce generation of test cases during refinement.
+Using problem descriptions as search queries can be difficult, because problems are generally intricate and include several components.
+Therefore, we propose an online search with query generation through planning.
+To address Challenge 2, we introduce generation of test cases during refinement.
 Several studies Chen et al.
@@
 It significantly simplifies the process of test case generation and facilitates its precise creation for complex code.
-To address the Challenge 3, this work prioritizes correctness testing in refinement.
-During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors.
-CoCoST incorporate both the execution output results and the errors into within the refinement prompts for LLMs to enhance the correctness.
-Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, Matplotlib charts.
-Thus, CoCoST propose serialization of input and output to convert them into understandable sequences before being processed by LLMs.
+To address Challenge 3, this work prioritizes correctness testing in refinement.
+During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking the existence of the errors.
+CoCoST incorporates both the execution output results and the errors within the refinement prompts for LLMs to enhance the correctness.
+Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, and Matplotlib charts.
+Thus, CoCoST proposes serialization of input and output to convert them into understandable sequences before being processed by LLMs.
 Particularly those are excessively long or non-textual modalities.
-We evaluate the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval).
-By comparing with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and average 9.47% on ClassEval.Moreover, we analyze and discovery that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem.
-In summary, our main contributions are:
-  • We propose the novel CoCoST framework to generate complex code, and can be automatic in real-world scenarios.
-  • We apply online search in code generation for the first time to our knowledge, in order to generate the complex code.
-  • We prioritize correctness testing in refinement with test case generation and serialization of input and output, in order to refine hidden bugs in complex code.
-  • We conduct experiments on DS-1000 and ClassEval datasets anddemonstrate the effectiveness and universality of CoCoST.
+We evaluated the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval).
+Compared with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and an average of 9.47% on ClassEval.Moreover, we analyze and discover that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem.
+In summary, our main contributions are as follows.
+  • We propose the novel CoCoST framework to generate complex code. CoCoST can be automatic in real-world scenarios.
+  • To generate complex code, we designed an online search method (query generation) in code generation for the first time to our knowledge.
+  • To refine hidden bugs in complex code, we prioritize correctness testing in refinement with test case generation and serialization of input and output data types.
+  • We conducted experiments on the DS-1000 and ClassEval datasets to demonstrate the effectiveness and universality of CoCoST.
 2 Related Work
@@
 Step 1: LLM is employed to strategize the problem and formulate queries based on the outlined steps.
-These queries enable the retrieval of diverse information from the internet.
+These queries enable the retrieval of various information from the internet.
 A high-quality initial code can be obtained through effective planning and leveraging internet information.
-Step 2: LLM generates test cases for testing the initial code.
-The test results serve as crucial inputs for the subsequent cycle of code refinement.
+Step 2: LLM generates test cases to test the correctness of the initial code.
+The serialization of test results serves as crucial input for the subsequent cycle of code refinement.
 Through iterative refinement processes, the quality of the initial code can be significantly improved.
@@
 In this work, we adopt a two-step approach for code generation, mirroring the way humans write code.
-The first step is retrieval, where relevant information is obtained through online search and utilized by LLMs to generate initial code.
+The first step is retrieval, where relevant information is obtained through an online search and utilized by LLMs to generate initial code.
 The second step is refinement, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
@@
 3.1.1 Generation Query through Planning
-In order to generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem.
+To generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem.
 The planning phase involves outlining the natural language steps $P=\{plan_{1},\ldots,plan_{N}\}$ required to address the problem.
@@
 3.1.2 Online Search
-For the above generated queries, we conduct an online search.
+For the above-generated queries, we conduct an online search.
 In this study, we use the online search API [^2] https://github.com/Nv7-GitHub/googlesearch for the search process as Equation (3).
@@
 Because their problems involve various logical operations, deriving answers directly without code-based computations is demanding.
-CoCoST introduces generation of test cases with LLMs to adapt to real-world scenarios.
+CoCoST introduces a generation of test cases with LLMs to adapt to real-world scenarios.
 $\left\{\begin{array}[]{ccl}\widehat{I}&\sim&p_{\theta}(.|D)\\ I&=&\{i_{1},\ldots,i_{N_{i}}\}\end{array}\right.$ (4)
@@
 4.1.1 Datasets
-We conduct experiences on two complex code generation datasets:
+We conduct experiences on two complex code-generation datasets:
 DS-1000 Lai et al.
@@
 Especially under the Diff-Rewrite perturbation setting, CoCoST exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CoCoST in generating complex code.
-CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refining the details to further enhance the correctness of the code.
+CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code and refining the details to further enhance the correctness of the code.
 For the ClassEval dataset, the results are shown in Table 2.
@@
 CoCoST significantly enhances the base model’s ability to generate complex code.
-Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
+Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and an average of 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
 Online search, generation of test cases, and serialization each contribute to the model’s performance improvements.
@@
 Online search outperforms offline retrieval in effectiveness and has a wider range of applicability.
-As shown in Table 1, using only online retrieval (the row w/o generation of test case) outperforms DocPrompting, which is an offline retrieval approach.
+As shown in Table 1, using only online retrieval (the row w/o generation of the test case) outperforms DocPrompting, which is an offline retrieval approach.
 Moreover, in real-world scenarios, as opposed to specific datasets, the types of problems encountered are more diverse.
@@
 During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important.
-After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval dataset.
+After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and an average of 2.7% and 0.3% on the ClassEval dataset.
 This indicates that the output result contributes more to the refinement process than the error.
@@
 Code generation necessitates models to have in-context learning abilities.
-The generated code should be built on all the above provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage.
+The generated code should be built on all the above-provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage.
 In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively.
@@
 This could be due to WizardCoder’s limited in-context learning ability, especially with the complex and lengthy prompts, hindering accurate context comprehension and code modification.
-4.5 Case Study and Error Analysis
+4.5 Cascade Analysis
+Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors.
+For the former, a discussion is provided in §4.5.1, while for the latter, an error analysis is conducted in §4.5.2.
+4.5.1 Analysis of Pipeline
+Regarding the generation of test cases, to prove that the generated test cases are comparable to the ground truth test cases, we substitute the generated test cases in CoCoST with ground truth test cases and conduct experiment on the DS-1000 dataset using GPT-4.
+The results showed that the performance was 68.70%, only marginally higher by 0.7% compared to the use of generated test cases (68.00%).
+This proves that the effects of both are comparable and that the generated test cases rarely lead to errors.
+Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval.
+On the DS-1000 dataset using GPT-4, this approach results in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%).
+This confirms that the generated plans are significantly effective.
+Figure 3: Case Study for Correctness Testing.
+4.5.2 Error Analysis
+For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements.
+Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect.
+For the former cases, it shows that CoCoST can fix some bad cases in the refinement stage even though the retrieval contents have some errors.
+For the later cases, we do observe some cases that are worthy of further research as potential directions for future work.
+First, the search content could be more detailed.
+E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but are exactly where the bug is in the code.
+Second, the search query could be more targeted.
+The descriptions of some queries are not specific enough in terms of some complicated problems.
+Therefore, more sub-queries are needed to help the model receive clearer instructions.
+Figure 4: Case Study for Online Retrieval.
+4.6 Case Study
 For the case study on online retrieval, refer to Figure 4.
@@
 The model refines the code based on the output, thus improving hidden errors and generating the correct code.
-Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors.
-For the former, a discussion is provided in §C.2, while for the latter, an error analysis is conducted in §C.3.
 5 Conclusion
-In conclusion, CoCoST introduces a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation.
+In this paper, we propose CoCoST, a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation.
 It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data.
@@
 Limitations
-The main limitation of our research is that it has underlying issues of exceeding the allowed times of accesses due to multiple calls to the Google Search API.
-Similarly, we also have made multiple API calls to test and enhance the performance of the GPT models.
-Ethics Policy
-This research does not pose any ethical concerns. The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
+The primary limitation of our research is exceeding the allowed access limits due to frequent calls to the Google Search API.
+Likewise, we made multiple API calls to test and optimize the performance of the GPT model.
+Ethics Statement
+The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
+All the datasets used in this paper have been reviewed to ensure that they do not contain personally identifiable information or offensive content.
+However, since these datasets are sourced from the Internet, potential bias may still be present.
+Furthermore, despite our careful review, the process involving the LLMs may inadvertently introduce inappropriate information into the evolved data.
+Acknowledgments
+We thank all anonymous reviewers for their valuable comments. Xinyi He and Zejian Yuan were supported in part by the National Key R&D Program of China (2023YFB4704900) and NSFC (61976170, 62088102).
 Appendix A Online Searching Detail
@@
 Without these test cases, they are entirely incapable of functioning.
+Table 5: Table of Main Results for different packages in DS-1000. All metric numbers are represented as percentages. The bold number indicates the highest performance.
+Method Pandas Numpy Matplotlib Tensorflow Scipy Sklearn Pytorch Total/Avg.
+CoCoST 59.45 75.91 75.48 71.11 61.32 63.48 77.94 68.00
+     + retrieve 51.89 70.91 68.39 66.67 52.83 70.43 60.29 62.10
+     + refine 55.67 72.73 74.19 64.44 54.72 60.00 70.59 64.10
+     GPT-4 only 52.23 70.45 67.74 55.56 50.00 64.35 55.88 60.20
 B.4 Ablation Study Details
@@
     Appendix C Experimantal Results
-    C.1 DS-1000 Results
     The main results for different packages in DS-1000 are shown in Table 5.
-    Table 5: Table of Main Results for different packages in DS-1000. All metric numbers are represented as percentages. The bold number indicates the highest performance.
-    Method Pandas Numpy Matplotlib Tensorflow Scipy Sklearn Pytorch Total/Avg.
-    CoCoST 59.45 75.91 75.48 71.11 61.32 63.48 77.94 68.00
-         + retrieve 51.89 70.91 68.39 66.67 52.83 70.43 60.29 62.10
-         + refine 55.67 72.73 74.19 64.44 54.72 60.00 70.59 64.10
-         GPT-4 only 52.23 70.45 67.74 55.56 50.00 64.35 55.88 60.20
     The results indicate that CoCoST shows a more pronounced effect on libraries whose inputs and outputs are more complex or more challenging for LLMs to intuitively understand, such as Matplotlib, TensorFlow, and PyTorch.
@@
 Consequently, CoCoST’s performance on Sklearn is not as strong as with the other libraries.
-    C.2 Analysis of Pipeline
-    Regarding the generation of test cases, to prove that the generated test cases are comparable to the ground truth test cases, we substitute the generated test cases in CoCoST with ground truth test cases and conduct experiment on the DS-1000 dataset using GPT-4.
-The results showed that the performance was 68.70%, only marginally higher by 0.7% compared to the use of generated test cases (68.00%).
-This proves that the effects of both are comparable and that the generated test cases rarely lead to errors.
-    Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval.
-On the DS-1000 dataset using GPT-4, this approach result in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%).
-This confirms that the generated plans are significantly effective.
-    C.3 Error Analysis
-    For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements.
-Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect.
-For the former cases, it shows that CoCoST is able to fix some bad cases in the refinement stage even though the retrieval contents have some errors.
-For the later cases, we do observe some cases that are worthy of further research as potential directions for future work.
-First, the search content could be more detailed.
-E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but is exactly where the bug in the code.
-Second, the search query could be more targeted.
-The descriptions of some queries are not specific enough in terms of some complicated problems.
-Therefore, more sub-queries are needed to help model receive clearer instructions.
-C.4 Case Study
-Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset.
-Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices.
-After unsuccessful attempts, the model turns to an online search through a self-generated query and integrates the research results as part of the next prompt.
-Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class.
-Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
-Figure 3: Case Study for Correctness Testing.
-Figure 4: Case Study for Online Retrieval.
 Appendix D Prompts of CoCoST